OpenStudy (anonymous):

If : \(\color{blue}{\text{2 + 6 + 10 + 14 + 18 + 22 + 26 + 30 + 34 + 38 = 200}}\), Then, \(\color{red}{\text{Add}}\) any \(\color{cyan}{\text{Five}}\) numbers and get \(\color{green}{100}\)..

OpenStudy (amistre64):

???

OpenStudy (amistre64):

methinks someones just practicing thier latex codeing

OpenStudy (mathmate):

You mean so that \( \color{blue}{\text{2 + 6 + 10 + 14 + 18 + 22 + 26 + 30 + 34 + 38}} \color{red}{\text{ + a + b + c + d + e }} \color{blue}{\text{=100}} \)

OpenStudy (mathmate):

Then any numbers such that a+b+c+d+e=-100 will do the job.

OpenStudy (anonymous):

This cannot be done as one can prove that one cannot take any 5 numbers, add them, and get 100: Any element is expressed as: 2 + 4(n - 1) Adding any 5 will give: (2)(5) + 4(x) = 100 where "x" is some integer. 4(x) = 90 x = 90/4 is not an integer. Cannot be done.